OpenStudy (anonymous):

evaluate each expression (-4) + (7-(-3)) x 6

OpenStudy (aroub):

=60x-4

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

this isn't an agebraic question is it? by x6 do you mean times 6?

OpenStudy (aroub):

oh i took it as an x

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

no its times

OpenStudy (anonymous):

10 *6=60 -4=56

OpenStudy (anonymous):

if i've read it right